Now, that you have a deeper understanding of the topic, I believe that you are ready to solve the problems below.


Mission Possible


1. The volume V of air remaining in an inflated balloon can be modeled by the function V = 1,000(0.85) where x represents the number of days that have passed since inflating the balloon. What is the reasonable domain for the situation? Explain.


2. The function f(x)=65,000(1.5) can be modeled the population of a city for x, the number of years that have passed since 2010. What inequality represents the reasonable range of the function based on the situation? Explain.


1
Expert's answer
2021-10-20T06:50:57-0400

1.

"V=1000(0.85)^x"

Since "x" represents the number of days that have passed since inflating the balloon, we assume that "x\geq0."

Since "V" represents the volume of air remaining in an inflated balloon, then "V\to0" as "x\to \infin."

Therefore the  reasonable domain for the situation is "[0, \infin)."


2.

"f(x)=65000(1.5)^x"

Since "x" represents the number of years that have passed since 2010, we assume that "x\geq0."


"f(0)=65000(1.5)^0=65000"

Since "1.5>1" the function "f(x)" increases for "x\geq0," "f(x)\to \infin" as "x\to \infin."

Therefore the  reasonable range of the function based on the situation is "[65000, \infin)."